# CLIP4STR: A Simple Baseline for Scene Text Recognition with Pre-trained Vision-Language Model

### v3

@@ III-C1 Decoding scheme
ViTSTR-S [12] ImageNet-1K Real(3.3M) 97.9 96.0 97.8 89.0 87.5 91.5 96.2 64.5 77.9
ABINet [11] From Scratch Real(3.3M) 98.6 98.2 98.0 90.5 88.7 94.1 97.2 72.2 85.0
PARSeqA [26] From Scratch Real(3.3M) 99.1 97.9 98.4 90.7 89.6 95.7 98.3 74.4 85.4
MAERec-B [63] Union14M-U Union14M-L 98.5 97.8 98.1 - 89.5 94.4 98.6 - -
CLIP4STR-B WIT 400M MJ+ST 97.7 95.2 96.1 87.6 84.2 91.3 95.5 79.8 87.0
CLIP4STR-L MJ+ST 98.0 95.2 96.9 87.7 84.5 93.3 95.1 82.7 88.8
CLIP4STR-B Real(3.3M) 99.2 98.3 98.3 91.4 90.6 97.2 99.3 77.5 87.5
CLIP4STR-L Real(3.3M) 99.5 98.5 98.5 91.3 90.8 97.4 99.0 79.8 89.2
@@
CLIP4STR-B RBU(6.5M) 99.5 98.3 98.6 91.4 91.1 98.0 99.0 79.3 88.8
CLIP4STR-L RBU(6.5M) 99.6 98.6 99.0 91.9 91.4 98.1 99.7 81.1 90.6
CLIP4STR-H DFN-5B [84] RBU(6.5M) 99.5 99.1 98.9 91.7 91.0 98.0 99.0 82.6 90.9
IV Experiment
IV-A Experimental Details
We instantiate CLIP4STR with CLIP-ViT-B/16, CLIP-ViT-L/14, and CLIP-ViT-H/14 [40]. Table II presents the main hyper-parameters of CLIP4STR. The reproducibility of CLIP4STR is verified by a third-party PyTorch implementation and recent works [64]. Please refer to Appendix A.
Test benchmarks  The evaluation benchmarks include IIIT5K [85], CUTE80 [86], Street View Text (SVT) [87], SVT-Perspective (SVTP) [88], ICDAR 2013 (IC13) [23], ICDAR 2015 (IC15) [24], and two occluded datasets – HOST and WOST [78]. Additionally, we utilize 3 recent large benchmarks: COCO-Text (9.8K samples; low-resolution, occluded text) [89], ArT (35.1K samples; curved and rotated text) [17], and Uber-Text (80.6K samples; vertical and rotated text) [18].
@@ IV-A Experimental Details
IV-B Comparison to State-of-the-art
We compare CLIP4STR with previous SOTA methods on 8 common STR benchmarks in Table III. CLIP4STR surpasses the previous methods by a significant margin, achieving new SOTA performance. Notably, CLIP4STR performs exceptionally well on irregular text datasets, such as IC15 (incidental scene text), SVTP (perspective scene text), CUTE (curved text line images), HOST (heavily occluded scene text), and WOST (weakly occluded scene text). This aligns with the examples shown in Figure 1&2 and supports our motivation for adapting CLIP as a scene text reader, as CLIP demonstrates robust identification of regular and irregular text. CLIP4STR exhibits excellent reading ability on occluded datasets, surpassing the previous SOTA by 7.8% and 5.4% in the best case on HOST and WOST, respectively. This ability can be attributed to the pre-trained text encoder and cross-modal decoder, which can infer missing characters using text semantics or visual features. The performance of CLIP4STR is also much better than CLIP-OCR [68] and CLIPTER [67], both of which work in a similar direction as CLIP4STR. This demonstrating that directly transferring CLIP into a STR reader is more effective than the distillation method [68] or utilizing CLIP features of the global image as auxiliary context [67].
In addition to the small-scale common benchmarks, we also evaluate CLIP4STR on three larger and more challenging benchmarks. These benchmarks primarily consist of irregular texts with various shapes, low-resolution images, rotation, etc. The results, shown in Table IV, further demonstrate the strong generalization ability of CLIP4STR. CLIP4STR outperforms the previous SOTA methods substantially on these three large and challenging benchmarks. At the same time, we observe that scaling CLIP4STR to 1B parameters does not bring much improvement in performance. CLIP4STR-L is comparable to CLIP4STR-H in most cases, while CLIP4STR-H is superior in recognizing occluded characters (WOST and HOST).
@@ V Empirical Study
This section presents our empirical study on adapting CLIP to STR. Without mention,the models are all trained on 3.3M real data, and IC15 dataset here contains 2,077 samples.
V-A Ablation Study of CLIP4STR
Table III&IV show that CLIP4STR achieves SOTA performance on 11 STR benchmarks. What are the sources of this high performance? We conduct ablation studies of different components in Table V, starting with the visual branch in Figure 3 as the baseline (accuracy 89.2%). The encoder is a ViT-S without pre-training. Then we apply the permuted sequence modeling (PSM) technique [26] to the visual decoder and follow the training recipe of PARSeq: 4$\times$8 patch size, the same learning rate for the encoder and decoder, and 20 training epochs. This brings a 0.7% improvement in accuracy. Next, we replace the encoder with the image encoder of CLIP-ViT-B/16. However, no significant gain is observed without adaptations. To unleash the potential of CLIP, we adjust the training recipe: using 16$\times$16 patch size, a small learning rate for CLIP encoders, a relatively large learning rate for decoders, and fewer training epochs — 16 (Section IV-A). The learning rate is searched automatically by Ray [101], and the best number of training epochs is decided by manual test. CLIP makes the model converge easier and faster, so the training recipe should change accordingly. To better grasp the contribution of PSM, we conducted an experiment where we removed PSM and achieved an accuracy of 90.0%. This emphasizes the significance of PSM.
Although the performance of visual branch is already very high (90.8%), the cross-modal branch further improves the accuracy by 0.4%, demonstrating its effectiveness. It is worth noting that utilizing CLIP features of all patches is crucial for character recognition. Only using the [CLASS] and [EOS] results in inferior performance – 90.6%. Moreover, the use of a large model — CLIP-ViT-L/14 further increases the accuracy by 0.7%. The large CLIP-ViT-L/14 converges faster than CLIP-ViT-B/16 for STR. It only requires 10 epochs of training on the Real(3.3M) data, much less than the training epochs of CLIP-ViT-B/16.
TABLE V: Ablation study of different components of CLIP4STR. Average accuracy on 9 benchmarks (14,315 samples) in Table III are reported. PSM is the permuted sequence modeling method proposed by PARSeq [26]. Recipe is the training recipe for CLIP4STR introduced in Section IV-A. Cross is the cross-modal branch. ✓  for [CLASS] means the decoders only use the [CLASS] and [EOS] of CLIP encoders rather than features of all tokens.
@@ V-B Parameter Freezing Options
V-C Comparison to Single-modality Pre-trained Model
In previous empirical studies, we see the effectiveness of CLIP as a STR backbone. Is VLM better than models pre-trained on single-modality data? To further clarify this question, Table VII presents the results of replacing the visual encoder in Figure 3 with a random initialized ViT, an ImageNet-1K [102] pre-trained ViT via DeiT [103] [^3] https://github.com/facebookresearch/deit, and an ImageNet-21K pre-trained ViT provided by Ridnik et al. [104]  [^4] https://github.com/Alibaba-MIIL/ImageNet21K. The training schedules including the learning rate and training epochs are kept the same as CLIP4STR. In Table VII, the ImageNet pre-trained models even perform worse than the model trained from scratch. Previous works also support this finding. PARSeq [26] trains its vision transformer from scratch rather than using a pre-trained model. TrOCR [14] uses pre-trained transformers from DeiT [103], BEiT [105], and RoBERTa [106], but it still post-pretrains them on 684M textlines from publicly available PDF files on the Internet. Table VII demonstrates the advantage of using a VLM learning under text supervision in scene text recognition.
Figure 5: CLIP-Adapter (left) and LST (right).
@@ V-D Parameter-efficient Adaptations
We test CLIP with two PEFT methods in this work, i.e., CLIP-Adapter [108] and Ladder Side-Tuning (LST) adapter [109]. Figure 5 shows the design of the two adapters. CLIP-Adapter adds two linear layers on the top of the frozen pre-trained VLM. We use the same architecture as the original implementation  [^5] https://github.com/gaopengcuhk/CLIP-Adapter and a residual addition ratio $\lambda=0.2$, which means that the original CLIP feature is multiplied by $0.8$. Ladder Side-Tuning (LST) uses a ladder side network as shown in Figure 5. We follow the original implementation  [^6] https://github.com/ylsung/Ladder-Side-Tuning and use the structure-pruned [110] CLIP model as the ladder side network. The CLIP features are downsampled by a factor of $1/r$ before entering the ladder side network to reduce the computation cost, and then upsampled by a factor of $r$ before output to match the original feature dimension. We also use the layer-dropping strategy in LST, which connects only the layers $[2,4,6,8,10,12]$ to the ladder side network, namely, the depth of LST is 6. This reduces the training cost.
@@
TABLE IX: Inference time of CLIP4STR. Average accuracy on 9 benchmarks (14,315 samples) in Table III are reported. AR stands for autoregressive decoding, and cloze stands for cloze-filling decoding manner (see Alogorithm 1). Iter. is the number of refinement iterations during decoding. Time is the average inference time per sample. Test on a single NVIDIA A100 40GB GPU.
@@
V-E Inference Time
Despite the good performance, adapting the pre-trained CLIP model introduces extra training and inference costs due to its large size. Table IX presents the inference time of CLIP4STR. The large transformer models slow down the inference speed of CLIP4STR. However, using a large ViT does not always improve accuracy, as Table VII shows, because of different pre-training strategies. The cross-modal branch also increases the inference time, but slightly (0.49ms), since the input sequence length of the text encoder is small (16, as explained in Sec. IV-A). Moreover, we can reduce the inference time of the cross-modal branch by replacing line 10~13 in Algorithm 1 with
Equation (9) uses the prediction of the visual branch as the input context instead of the previous prediction of the cross-modal branch, avoiding repeated runs of the cross-modal decoder. However, this slightly decreases the performance. The ViT-L backbone also increases the inference time. Clearly, for CLIP4STR, there is a trade-off between recognition accuracy and inference speed. Besides, Table IX also shows that more iterative refinement times (a large $T_{i}$ at line 14 in Algorithm 1) will not bring further improvement in accuracy, so we just set $T_{i}=1$ in practice.
V-F Qualitative results
Figure 6: Qualitative results of CLIP4STR-B.
Figure 6 shows a few qualitative results of CLIP4STR on IC15 (incidental scene text), SVTP (perspective scene text), CUTE (curved text line images), and HOST (heavily occluded scene text). CLIP4STR can robustly read scene text that is curved, occluded, blurred, or rotated. This matches its state-of-the-art performance in Table III and verifies our motivation for adapting CLIP to STR in Section I.
@@ V-G Results on Cleaned Benchmarks
Recently, Yang et al[100] correct the ground truth of mislabeled samples and present cleaned versions of IIIT5K, SVT, IC13, IC15, SVTP, and CUTE  [^7] https://github.com/Xiaomeng-Yang/STR_benchmark_cleansed. Table X shows the results of CLIP4STR on these cleaned benchmarks. CLIP4STR still achieves SOTA performance on these cleaned benchmarks.
@@ VI Conclusion
We present CLIP4STR, a method that leverages CLIP for STR. It has a dual encoder-decoder architecture: a visual branch for initial prediction and a cross-modal branch for refinement. CLIP4STR achieves state-of-the-art results on 11 STR benchmarks, showing that CLIP is a powerful scene text reader and that vision-language pre-training is beneficial for STR. We also conduct a comprehensive empirical study to explain how CLIP adapts to STR. We hope that CLIP4STR can serve as a simple but strong baseline for future STR research with VLMs.
Appendix A Reproducibility of CLIP4STR
A thirty-party open-sourced code at https://github.com/VamosC/CLIP4STR reproduces the performance of CLIP4STR. A recent work [64] also demonstrate the reproducibility of CLIP4STR.
Appendix B Detail Explanation of the Inference Process
Here we provide an explanation of the inference process in Algorithm 1. Given an image $\bm{x}$, the initial step involves obtaining the image feature $\bm{F}_{i}\leftarrow h(\bm{x})$. This image feature $\bm{F}_{i}$ is then forwarded to the visual decoder to generate the visual prediction $\bm{y}^{i}$, with the blank context (denoted as token [B]) serving as the initial condition (line 1). Subsequently, the visual decoder operates in an autoregressive manner, utilizing previous predictions as context for subsequent ones (lines 4-7). Once the prediction is obtained from the visual branch, the linguistic feature is derived by inputting the visual prediction $\bm{y}^{i}$ along with the cross-model feature $\bm{F}_{c}$ into the text encoder (line 8). Similar to the decoding process of the visual decoder, the cross-modal decoder generates predictions $\bm{y}$ in an autoregressive fashion (lines 10–13). Upon acquiring predictions $\bm{y}^{i}$ and $\bm{y}$, they are employed to update the context $\bm{c}$ during the refinement process (lines 15 and 18). Notably, while the decoder previously produced $\bm{y}^{i}$ and $\bm{y}$ in an autoregressive manner, a different approach is adopted in lines 14–20, where a cloze mask is utilized. This entails providing information about other characters in the word as context when predicting a certain character. For further insights into the workings of autoregressive and cloze masks, please refer to Table I.

### v4

@@ IV-A Experimental Details
ViTSTR-S [10] ImageNet-1K Real(3.3M) 97.9 96.0 97.8 89.0 87.5 91.5 96.2 64.5 77.9 64.2
ABINet [9] From Scratch Real(3.3M) 98.6 98.2 98.0 90.5 88.7 94.1 97.2 72.2 85.0 70.1
PARSeqA [24] From Scratch Real(3.3M) 99.1 97.9 98.4 90.7 89.6 95.7 98.3 74.4 85.4 73.1
MAERec-B [61] Union14M-U Union14M-L 98.5 97.8 98.1 89.5 94.4 98.6
CLIP4STR-B WIT 400M MJ+ST 97.7 95.2 96.1 87.6 84.2 91.3 95.5 79.8 87.0 57.1
CLIP4STR-L MJ+ST 98.0 95.2 96.9 87.7 84.5 93.3 95.1 82.7 88.8 59.2
CLIP4STR-B Real(3.3M) 99.2 98.3 98.3 91.4 90.6 97.2 99.3 77.5 87.5 81.8
CLIP4STR-L Real(3.3M) 99.5 98.5 98.5 91.3 90.8 97.4 99.0 79.8 89.2 84.9
@@
CLIP4STR-B RBU(6.5M) 99.5 98.3 98.6 91.4 91.1 98.0 99.0 79.3 88.8 83.5
CLIP4STR-L RBU(6.5M) 99.6 98.6 99.0 91.9 91.4 98.1 99.7 81.1 90.6 85.9
CLIP4STR-H DFN-5B [96] RBU(6.5M) 99.5 99.1 98.9 91.7 91.0 98.0 99.0 82.6 90.9 86.5
@@
IV-B Comparison to State-of-the-art
We compare CLIP4STR with previous SOTA methods on 10 common STR benchmarks in Table III. CLIP4STR surpasses the previous methods by a significant margin, achieving new SOTA performance. Notably, CLIP4STR performs exceptionally well on irregular text datasets, such as IC15 (incidental scene text), SVTP (perspective scene text), CUTE (curved text line images), HOST (heavily occluded scene text), and WOST (weakly occluded scene text). This aligns with the examples shown in Fig. 1&2 and supports our motivation for adapting CLIP as a scene text reader, as CLIP demonstrates robust identification of regular and irregular text. CLIP4STR exhibits excellent reading ability on occluded datasets, surpassing the previous SOTA by 7.8% and 5.4% in the best case on HOST and WOST, respectively. This ability can be attributed to the pre-trained text encoder and cross-modal decoder, which can infer missing characters using text semantics or visual features. The performance of CLIP4STR is also much better than CLIP-OCR [66] and CLIPTER [65], both of which work in a similar direction as CLIP4STR. This demonstrates that directly transferring CLIP into a STR reader is more effective than the distillation method [66] or utilizing CLIP features of the global image as auxiliary context [65].
@@ IV-B Comparison to State-of-the-art
In addition to the small-scale common benchmarks, we also evaluate CLIP4STR on 3 larger and more challenging benchmarks. These benchmarks primarily consist of irregular texts with various shapes, low-resolution images, rotation, etc. The results, shown in Table IV, further demonstrate the strong generalization ability of CLIP4STR. CLIP4STR substantially outperforms the previous SOTA methods on these three large and challenging benchmarks. At the same time, we observe that scaling CLIP4STR to 1B parameters does not bring much improvement in performance. CLIP4STR-L is comparable to CLIP4STR-H in most cases, while CLIP4STR-H is superior in recognizing occluded characters (WOST, HOST, OCTT).
@@ V Empirical Study
This section presents our empirical study on adapting CLIP to STR. Without mention, the models are all trained on 3.3M real data, and the IC15 dataset here contains 2,077 samples. The average accuracy reported in this section is calculated over the first 9 benchmarks (14,315 samples) in Table III.
TABLE V: Ablation study of different components of CLIP4STR. PSM is short for the permuted sequence modeling technique [24]. Recipe represents the training recipe for CLIP4STR in §IV-A. Cross denotes the cross-modal branch. [CLASS] with a ✓mark means the decoders only use the [CLASS] and [EOS] of CLIP encoders rather than features of all tokens (refer to §III-A).
@@
V-A Ablation Study of CLIP4STR
Table III&IV show that CLIP4STR achieves SOTA performance on 11 STR benchmarks. What are the sources of this high performance? We conduct ablation studies of different components in Table V, starting with the visual branch in Fig. 3 as the baseline (accuracy 89.2%). The encoder is a ViT-S without pre-training. Then we apply the permuted sequence modeling (PSM) technique [24] to the visual decoder and follow the training recipe of PARSeq: 4$\times$8 patch size, the same learning rate for the encoder and decoder, and 20 training epochs. This brings a 0.7% improvement in accuracy. Next, we replace the encoder with the image encoder of CLIP-ViT-B/16. However, no significant gain is observed without adaptations. To unleash the potential of CLIP, we adjust the training recipe: using 16$\times$16 patch size, a small learning rate for CLIP encoders, a relatively large learning rate for decoders, and fewer training epochs — 16 (§IV-A). The learning rate is searched automatically by Ray [99], and the best number of training epochs is decided by manual test. CLIP makes the model converge easier and faster, so the training recipe should change accordingly. To better grasp the contribution of PSM, we conducted an experiment where we removed PSM and achieved an accuracy of 90.0%.
Although the performance of the visual branch is already very high (90.8%), the cross-modal branch further improves the accuracy by 0.4%, demonstrating its effectiveness. It is worth noting that utilizing CLIP features of all patches is crucial for character recognition. Only using the [CLASS] and [EOS] results in inferior performance – 90.6%. Moreover, the use of a large model — CLIP-ViT-L/14 further increases the accuracy by 0.7%. The large CLIP-ViT-L/14 converges faster than CLIP-ViT-B/16 for STR. It only requires 10 epochs of training on the Real(3.3M) data, much less than the training epochs of CLIP-ViT-B/16.
@@ V-B Parameter Freezing Options
V-C Comparison to Single-modality Pre-trained Model
In previous empirical studies, we see the effectiveness of CLIP as a STR backbone. Is VLM better than models pre-trained on single-modality data? To further clarify this question, Table VII presents the results of replacing the visual encoder in Fig. 3 with a random initialized ViT, an ImageNet-1K [100] pre-trained ViT via DeiT [101] [^3] https://github.com/facebookresearch/deit, and an ImageNet-21K pre-trained ViT provided by Ridnik et al. [102]. The training schedules including the learning rate and training epochs are kept the same as CLIP4STR. In Table VII, the ImageNet pre-trained models even perform worse than the model trained from scratch. Previous works also support this finding. PARSeq [24] trains its vision transformer from scratch rather than using a pre-trained model. TrOCR [12] uses pre-trained transformers from DeiT [101], BEiT [103], and RoBERTa [104], but it still post-pretrains them on 684M textlines from publicly available PDF files on the Internet.
ImageNet classification pre-training does not align well with STR. Classifying objects in an image does not help the model learn specific information about the text within the image. For example, two images – one of a cat and one of a dog – both containing the text “park” cause the model to learn contradictory information about the same text. In contrast, the vision encoders in CLIP can accurately perceive text signals due to the presence of multi-modal neurons [20] (§II-A), making CLIP a strong backbone for STR.
@@ V-D Parameter-efficient Adaptations
Figure 5: CLIP-Adapter (left) and LST (right).
We test CLIP with two PEFT methods in this work, i.e., CLIP-Adapter [106] and Ladder Side-Tuning (LST) adapter [107]. Fig. 5 shows the design of the two adapters. CLIP-Adapter adds two linear layers on the top of the frozen pre-trained VLM. We use the same architecture as [106]: a residual addition ratio $\lambda=0.2$, which means that the original CLIP feature is multiplied by $0.8$. Ladder Side-Tuning (LST) uses a ladder side network as shown in Fig. 5. Following [107], we use the structure-pruned [108] CLIP model as the ladder side network. The CLIP features are downsampled by a factor of $1/r$ before entering the ladder side network to reduce the computation cost, and then upsampled by a factor of $r$ before output to match the original feature dimension. We also use the layer-dropping strategy in LST, which connects only the layers $[2,4,6,8,10,12]$ to the ladder side network, namely, the depth of LST is 6. This reduces the training cost.
@@
V-E Inference Time
Despite the good performance, adapting the pre-trained CLIP model introduces extra training and inference costs due to its large size. Table IX presents the inference time of CLIP4STR. The large transformer models slow down the inference speed of CLIP4STR. However, using a large ViT does not always improve accuracy, as Table VII shows, because of different pre-training strategies. The cross-modal branch also increases the inference time, but slightly (0.49ms), since the input sequence length of the text encoder is small (16, as explained in §IV-A). Moreover, we can reduce the inference time of the cross-modal branch by replacing line 10$\sim$13 in Alg. 1 with
Eq. (8) uses the prediction of the visual branch as the input context instead of the previous prediction of the cross-modal branch, avoiding repeated runs of the cross-modal decoder. However, this slightly decreases the performance. The ViT-L backbone also increases the inference time. Clearly, for CLIP4STR, there is a trade-off between recognition accuracy and inference speed. Besides, Table IX also shows that more iterative refinement times (a large $T_{i}$ at line 14 in Alg. 1) will not bring further improvement in accuracy, so we just set $T_{i}\!=\!1$ in practice.
TABLE IX: Inference time of CLIP4STR. AR stands for autoregressive decoding, and cloze stands for cloze-filling decoding manner (refer to Table I). Iter. is the number of refinement steps during decoding. Time is the average inference time per sample on a single NVIDIA A100 40GB.
@@ V-E Inference Time
V-F Qualitative results
Figure 6: Qualitative results of CLIP4STR-B.
Fig. 6 shows qualitative results of CLIP4STR on IC15 (incidental text), SVTP (perspective text), CUTE (curved text), and HOST (heavily occluded). CLIP4STR can robustly read scene text that is curved, occluded, blurred, or rotated, showing its great robustness. Meanwhile, we find that CLIP4STR has a strong ability to complement incomplete characters. In the last two cases in Fig. 6, CLIP4STR predicts an additional “n” character. This capability may stem from the semantic understanding of the pre-trained CLIP model. However, the accuracy of this complement is uncertain, and we currently cannot control this behavior in CLIP4STR.
@@ V-G Results on Cleaned Benchmarks
Recently, Yang et al[98] correct the ground truth of mislabeled samples and present cleaned versions of IIIT5K, SVT, IC13, IC15, SVTP, and CUTE. Table X shows the results of CLIP4STR on these cleaned benchmarks. CLIP4STR still achieves SOTA performance on these cleaned benchmarks.
@@ VI Conclusion
We present CLIP4STR, a method that leverages CLIP for STR. It has a dual encoder-decoder architecture: a visual branch for initial prediction and a cross-modal branch for refinement. CLIP4STR achieves state-of-the-art results on 13 STR benchmarks, showing that CLIP is a powerful scene text reader and that vision-language pre-training benefits STR. We also conduct a comprehensive empirical study to explain how CLIP adapts to STR. We hope CLIP4STR can serve as a simple but strong baseline for future STR research with VLMs.
Appendix A Detail Explanation of the Inference Process
Here we provide an explanation of the inference process in Alg. 1. Given an image $\bm{x}$, the initial step involves obtaining the image feature $\bm{F}_{i}\leftarrow h(\bm{x})$. This image feature $\bm{F}_{i}$ is then forwarded to the visual decoder to generate the visual prediction $\bm{y}^{i}$, with the blank context (denoted as token [B]) serving as the initial condition (line 1). Subsequently, the visual decoder operates in an autoregressive manner, utilizing previous predictions as context for subsequent ones (lines 4-7). Once the prediction is obtained from the visual branch, the linguistic feature is derived by inputting the visual prediction $\bm{y}^{i}$ along with the cross-model feature $\bm{F}_{c}$ into the text encoder (line 8). Similar to the decoding process of the visual decoder, the cross-modal decoder generates predictions $\bm{y}$ in an autoregressive fashion (lines 10–13). Upon acquiring predictions $\bm{y}^{i}$ and $\bm{y}$, they are employed to update the context $\bm{c}$ during the refinement process (lines 15 and 18). Notably, while the decoder previously produced $\bm{y}^{i}$ and $\bm{y}$ in an autoregressive manner, a different approach is adopted in lines 14–20, where a cloze mask is utilized. This entails providing information about other characters in the word as context when predicting a certain character. For further insights into the workings of autoregressive and cloze masks, please refer to Table I.